### v1

# Quantum hash function using discrete-time quantum walk on Hanoi network

Abstract
Quantum walk based hash functions have attracted a lot of attention in recent years because of its faster execution time and robust resistance against attacks compared to classical hash functions. It has been observed that the underlying graph and the way message controls the quantum walk iteration steps play a crucial role for the robustness of the hash function. We propose a quantum hash function based on the discrete-time quantum walk on a Hanoi network–a one dimensional periodic lattice with extra long-range edges of a specific form–which is highly collision resistant. The message bits of our scheme control the flow of probability amplitude through the extra long-range edges and the conditional shift operators. Our method even works for messages with small bit-lengths, contrary to most of the quantum walk based hash functions defined on a cycle, which usually work for messages with bit-lengths more than the length of the cycle.
Quantum Hash function; Quantum walk; Hanoi network; Collision resistance
I Introduction
Hash function [1]–a mathematical function–maps an arbitrary length message, $m$, to a fixed length number h(m)h(m)–known as the hash value. The process of generating hash function is such that, given a message, $m$, it is easy to obtain its corresponding hash value, $h(m)$, but from a hash value it is very difficult to recover the original message. Hash function has several important applications, including message authentication, public key infrastructure, and digital signature. There exist several classical hash functions, such as SHA-1, SHA-256, SHA-512, and MD5. However, some of them are vulnerable to attacks and suffer from security issues, which lead to the search beyond the realm of classical algorithm.
Quantum computing [2], on the other hand, offers a great opportunity to develop algorithms, which are fast and secure. For example, Grover’s algorithm [3, 4, 5] to search unsorted database quadratically faster than classical algorithm and Shor’s [6, 7] prime number factorization algorithm running in polynomial time are some of the well known quantum algorithms to mention. Quantum hash functions [8, 9, 10, 11, 12, 13, 14] offer unparalleled security based on the properties of quantum mechanics such as superposition, and entanglement. The basic element of a quantum algorithm is the initial quantum state, which is an element of the Hilbert space. In the case of quantum hash function, quantum states, obtained from different messages, need to be well separated in oder to avoid collision. It also obey other properties of the hash functions, which include easy to obtain hash value, deterministic nature of the generated hash values, hard to obtain message from the hash value etc.
Quantum walk [15]–a universal quantum computing primitive–is a suitable option for many quantum algorithms, such as quantum walk search [16, 17, 18] with some of the latest works reported in refs. [19, 20, 21], element distinctness [22] and many more including edge detection of digital images [23].
Recently, discrete-time quantum walk based schemes for the hash function have attracted a lot of attention [24, 25, 26, 29, 30, 31, 32, 33, 34, 35, 36]. Mostly a one-dimensional periodic lattice is involved for these schemes. For example, in ref. [24], quantum walk by two interacting particles on a one-dimensional periodic lattice, controlled by the message, has been exploited to obtain hash function. The coin operator is controlled by the message bits, i.e., depending on whether the bit value is 0 or $1$, the coin operator of the corresponding iteration involves $I$ interaction or $\pi$-phase interaction respectively. In this work, the authors considered two $2\times 2$ coin operators for the two coin spaces of the system. However, instead of two separate coin operators of the previous example, in refs. [25, 26] a $4\times 4$ coin operator is used for both the coin spaces, while the quantum walk is performed on a one-dimensional lattice. Also quantum hash function based on continuous-time quantum walk has been studied [27, 28]. So far, quantum walk based hash function has been studied in diverse underlying graphs, which include one-dimension lattice [29], two-dimensional periodic lattice with alternate quantum walk [30, 31] and Johnson graph of $J(n,1)$ [32], which is a complete graph.
While implementing hash function using quantum walk, there are many aspects which can influence the property of the function, which include, the graph on which quantum walk is performed, initial state of the walk, the way message bits control the evolution operator, and post processing of measured final quantum state. Usually, only the coin operator of the evolution operator is controlled by the message bits.
In the present article, we propose a quantum walk based hash function, where the one-dimensional periodic lattice is attached with a special form of long-range edges. Specifically, we use the Hanoi network of degree four (HN4) [37]. Note that, both the Hanoi networks of degree three (HN3) and four (HN4) have been used to study quantum walk search in the literature [38, 39]. Because of these long-range edges, we can efficiently obtain hash values for a message of small length, whose number of bits is smaller than the number of vertices of the graph for the quantum walk. On the other hand, most of the quantum walk based hash functions with a one-dimensional periodic lattice work only for the messages with bit-length more than the number of vertices in the graph. Also, both the coin and the shift operators in our method are controlled by the message bits. It allows us to get a hash function—highly resistant to collision.
This article is organized in the following fashion: A brief description of the quantum walk on HN4 network is presented in section II. In section III, a method for generating the hash function using this network is presented. In section IV statistical performance of our hash function, described in the previous section, has been analyzed. Under this section, sensitivity of the hash values to small changes in the messages, diffusion and confusion property, uniform distribution property, collision analysis and resistance to birthday attack are discussed. In section V a comparison of our quantum hash function with other quantum hash functions is done. Finally we conclude in section VI with a discussion.
II Quantum walk on Hanoi network
Hanoi network (HN) is generated from the sequence of numbered disks of the Tower-of-Hanoi problem [40]. Based on the number of edges at each vertex, there exist two important type of networks [37]: Hanoi network of degree three (HN3) and of degree four (HN4). Extensions of these typical Hanoi networks with even more degrees exist, but in this article, we only consider HN4 network for our study of the hash function. The main structure of this network is a one-dimensional periodic lattice of $N_{v}=2^{n}$ vertices. Like the one-dimensional lattice, it has two regular edges at each vertex $1\leq x_{v}\leq 2^{n}$, which are connected to its two nearest neighbor vertices. Additionally, it also has long-range edges, which give the small-world structure to the Hanoi network.
Fig. 1 represents a HN4 network with $N_{v}=16$ vertices. The main structure of the network is the solid black colored cycle with sixteen vertices represented by sixteen stars. Long-range edges are represented by dashed black color curves. Vertices $x_{v}$ can be written in terms of the parameters $i$ and $j$ as
$x_{v}=2^{i}\left(2j+1\right)\,$, (1)
Note that $0\leq i\leq n$ and $0\leq j\leq j_{max}=\lfloor 2^{n-i-1}-1/2\rfloor$ uniquely identify the vertices $x_{v}$. Alternatively, $N$ basis states $|x_{v}\rangle$ of the network can be expressed as $|x_{v}\rangle=|i,j\rangle$.
Figure 1: HN4 network with $N=2^{4}=16$ vertices. Solid and dashed black curves are regular and long-range edges respectively.
There exist four edges of the HN4 network: two regular edges, which come from the one-dimensional periodic lattice and two extra long-range edges connecting $|i,j\rangle$ to $|i,j-1\rangle$ and $|i,j+1\rangle$ respectively, for $i\neq n-1,n$. Also, there are two self-loops at $|n-1,0\rangle$ and $|n,0\rangle$ respectively. In fig. 1, we can see each vertex has two solid black colored regular edges and two dashed black colored long-range edges. Two of the long-range edges form undirected self-loops at $|n-1=3,0\rangle$ and $|n=4,0\rangle$ respectively.
The Hilbert space $\mathcal{H}=\mathcal{H}_{C}\times\mathcal{H}_{V}$ of this network is a tensor product of the coin space $\mathcal{H}_{C}$ and the vertex space $\mathcal{H}_{V}$ respectively. Quantum walk in this space is obtained by the coin operator, $C$, which acts on a state of the $4$-dimensional coin space, followed by the shift operation, $S$, which acts on the combined coin and $N_{v}$-dimensional vertex spaces. The initial state for the quantum walk is of the form
$|\psi_{in}\rangle=|\psi_{c}\rangle\otimes|\psi_{v}\rangle\,$, (2)
where $|\psi_{c}\rangle$ and $|\psi_{v}\rangle$ are the normalized state of the coin and vertex space respectively. Specific form of these states depend our choice, so that we can generate highly secured hash function.
The evolution operator for the quantum walk is given by
$\mathcal{U}=S\left(C\otimes I\right)\,$, (3)
which acts repeatedly on the initial state $|\psi_{in}\rangle$ to let the state evolve. However, in the case of hash function generation we need to modify the evolution operator $\mathcal{U}$ in each iteration step, which will be discussed in detail in the next section.
One of the important properties of the quantum walk on a finite one-dimensional lattice with $t$ time steps is that, for even $N$, probability distribution at location $x$ is zero [15] when $x+t$ is odd as can be seen from fig. 2(a). Therefore, the probability distribution is zero at even or add vertices for odd or even $t$ respectively. However, for odd $N$ the probability distribution is nonzero at all vertex locations for large enough $t$. So, normally, one cannot use even length one-dimensional periodic lattice for the generation of hash function, because it leads to very high collision rates due the predictable periodic zero probability distributions. The HN4 network, which we use in this study, allows us to consider even-length one-dimensional periodic lattice with no periodic zero-probability distribution as can be seen from fig. 2(b). Also, because of its two long-range edges we can manipulate probability distribution through the coin operator in such a way that the generated hash function has very high collision resistance. Note that, lively quantum walk with even lively-parameter studied in ref. [31] also does not have periodic zero probability distributions.
Figure 2: Probability distribution of a pair of randomly generated 99-bit length messages (message 11, message 22) as a function of the vertex location for the quantum walk on (a) a one-dimensional periodic lattice and (b) on a HN4 network for $N=16$, $t=9$.
Messages/hash values
1~~1 $m_{1}$ $111110010011000$
$h(m_{1})$ ‘6EEA’ ‘6CA5’ ‘3C9E’ ‘7333’ ‘5A78’ ‘3598’ ‘4AFC’ ‘A19D’ ‘5455’ ‘6DE2’ ‘3646’ ‘4518’ ‘4F38’ ‘7CC8’ ‘5CCF’ ‘6AAC’
2~~2 $m_{2}$ $111110000011000$
$h(m_{2})$ ‘858C’ ‘7492’ ‘7CB8’ ‘4377’ ‘3CBE’ ‘1F10’ ‘6794’ ‘1EBB’ ‘8C5B’ ‘46A6’ ‘3407’ ‘5F21’ ‘410F’ ‘8429’ ‘4592’ ‘937B’
3~~3 $m_{3}$ $111110010111000$
$h(m_{3})$ ‘5D89’ ‘17D0’ ‘346E’ ‘4933’ ‘5818’ ‘5D48’ ‘6255’ ‘53FB’ ‘359F’ ‘68BF’ ‘5053’ ‘73FD’ ‘84A2’ ‘3D01’ ‘9E02’ ‘942D’
4~~4 $m_{4}$ $1111110010011000$
$h(m_{4})$ ‘8414’ ‘9AF6’ ‘423D’ ‘500E’ ‘5A2A’ ‘49AD’ ‘79AC’ ‘7E39’ ‘3D09’ ‘39F3’ ‘6B34’ ‘5509’ ‘4047’ ‘61E0’ ‘5511’ ‘6065’
5~~5 $m_{5}$ $11111010011000$
$h(m_{5})$ ‘3546’ ‘2B9D’ ‘98E3’ ‘2DDF’ ‘6425’ ‘3B59’ ‘6351’ ‘69EF’ ‘87F7’ ‘4B93’ ‘A763’ ‘56F7’ ‘58E2’ ‘17C0’ ‘6ECF’ ‘54F1’
Table 1: Messages in binary forms and corresponding hash values in hexadecimal forms.
III Hash function with quantum walk on Hanoi network
To exploit quantum walk for the generation of hash function, the evolution operator $\mathcal{U}$ needs to be controlled by a message. Usually, only the coin operator of the evolution operator is controlled by the message bits. However, in our approach we control both the coin and shift operator, based on the bit-values(0/1)(0/1) of a message, as follows:
Bit-value =0=0: The evolution operator used in this case is the following:
$\mathcal{U}_{0}=S_{0}\left(C_{0}\otimes I\right)\,$, (4)
where we use the following Grover operator for the coin
$C_{0}=\left(2|\psi_{0}\rangle\langle\psi_{0}|-I\right)\,$, (5)
with
$|\psi_{0}\rangle=\mathcal{N}_{0}\left(|0\rangle+|1\rangle+\sqrt{l_{00}}|2\rangle+\sqrt{\tilde{l}_{00}}|3\rangle\right)\,$,
and $\mathcal{N}_{0}=1/(2+l_{00}+\tilde{l}_{00})$. The flip-flop shift operator is composed of three parts
$S_{0}=S_{00}+S_{01}+S_{02}\,.$ (6)
Note that, flip-flop shift operator is essential in spatial search on graph by discrete-time quantum walk [16, 18]. The first part of eq. (6)
$S_{00}=\sum^{N_{v}}_{x_{v}=1}\left(|1\rangle\langle 0|\otimes|x_{v}+1\rangle\langle x_{v}|\hskip 71.13188pt\right$.
$\left.+|0\rangle\langle 1|\otimes|x_{v}-1\rangle\langle x_{v}|\right)\,$,
is the usual flip-flop shift operator associated with the regular edges $|0\rangle$ and $|1\rangle$ of the HN4 network. The second part
$S_{01}=\sum^{n-2}_{i=0}\sum^{j_{max}}_{j=0}\left(|3\rangle\langle 2|\otimes|i,j+1\rangle\langle i,j|\hskip 56.9055pt\right$.
$\left.+|2\rangle\langle 3|\otimes|i,j-1\rangle\langle i,j|\right)\,$,
is the flip-flop shift operator associated with the long-range edges $|2\rangle$ and $|3\rangle$, and the third part
$S_{02}=\left(|2\rangle\langle 2|+|3\rangle\langle 3|\right)\otimes\hskip 113.81102pt$
$\left(|n-1,0\rangle\langle n-1,0|+|n,0\rangle\langle n,0|\right)\,$,
is associated with the two self-loops formed by the long-range edges at (i=n1,0)(i=n-1,0), and $(i=n,0)$ respectively. Note that, the four-dimensional coin space is formed by two regular edges $|0\rangle,|1\rangle$ and two long-range edges $|2\rangle,|3\rangle$.
Bit-value =1=1: The evolution operator in this case is the following:
$\mathcal{U}_{1}=S_{1}\left(C_{1}\otimes I\right)\,$, (7)
where we use Grover operator of the form
$C_{1}=\left(2|\psi_{1}\rangle\langle\psi_{1}|-I\right)\,$, (8)
with
$|\psi_{1}\rangle=\mathcal{N}_{1}\left(|0\rangle+|1\rangle+\sqrt{l_{11}}|2\rangle+\sqrt{\tilde{l}_{11}}|3\rangle\right)\,$,
and $\mathcal{N}_{1}=1/(2+l_{11}+\tilde{l}_{11})$. The regular shift operator, here, like the previous case, is composed of three parts
$S_{1}=S_{10}+S_{11}+S_{12}$ (9)
The first part
$S_{10}=\sum^{N_{v}}_{x_{v}=1}\left(|0\rangle\langle 0|\otimes|x_{v}+1\rangle\langle x_{v}|\hskip 71.13188pt\right$.
$\left.+|1\rangle\langle 1|\otimes|x_{v}-1\rangle\langle x_{v}|\right)\,$,
second part
$S_{11}=\sum^{n-2}_{i=0}\sum^{j_{max}}_{j=0}\left(|2\rangle\langle 2|\otimes|i,j+1\rangle\langle i,j|\hskip 56.9055pt\right$.
$\left.+|3\rangle\langle 3|\otimes|i,j-1\rangle\langle i,j|\right)\,$,
and the third part
$S_{12}=\left(|2\rangle\langle 2|+|3\rangle\langle 3|\right)\otimes\hskip 113.81102pt$
$\left(|n-1,0\rangle\langle n-1,0|+|n,0\rangle\langle n,0|\right)\,.$
The final state for the hash function for a general message $m=`b_{s}\cdots b_{2}b_{1}b_{0}^{\prime}$ in binary form with a bit-length $s$ is of the form
$|\psi_{f}\rangle=\mathcal{U}_{b_{0}}\mathcal{U}_{b_{1}}\cdots\mathcal{U}_{b_{s}}|\psi_{in}\rangle\,.$ (10)
For example, for a message $m=`110100^{\prime}$, the final state would be $|\psi_{f}\rangle=\mathcal{U}_{0}\mathcal{U}_{0}\mathcal{U}_{1}\mathcal{U}_{0}\mathcal{U}_{1}\mathcal{U}_{1}|\psi_{in}\rangle$. The iteration steps for the generation of our hash function, obtained from eq. (10) increases linearly with the bit-length $s$ of the message. Although, the results presented in this article are based on noiseless analysis, in real-life scenarios, a long bit-length can transform a quantum probability distribution of the quantum walk into a classical Gaussian-like distribution in the presence of unitary noise [41]. Correction methods such as error correction [42], and decoherence free sub-spaces [43] require significant overhead, which is a luxury in the current NISQ era. The effect of noise can be minimized by reducing the number of iteration steps of our scheme significantly by adopting an approach from ref. [33], where each evolution operator is controlled by two bits of the message instead of one in eq. (10), without compromising the performance our hash function. The final state for the hash function then becomes
$|\psi_{f}\rangle=\mathcal{U}_{b_{0}b_{1}}\mathcal{U}_{b_{2}b_{3}}\cdots\mathcal{U}_{b_{s-1}b_{s}}|\psi_{in}\rangle\,$, (11)
where $\mathcal{U}_{ij}=S_{i}\left(C_{ij}\otimes I\right)$. Four different coin operators $C_{00},C_{01},C_{10},C_{11}$, needed for this method, can be obtained by appropriately choosing two parameters associated with the long-range edges in the underlying coin state. For example, for a message $m=`110100^{\prime}$, the final state would now be $|\psi_{f}\rangle=\mathcal{U}_{00}\mathcal{U}_{10}\mathcal{U}_{11}|\psi_{in}\rangle$. The required iteration steps for the final state $|\psi_{f}\rangle$ can further be reduced by adopting other similar methods presented in ref. [33].
The final probability for the walker to be at $x_{v}$ location can be evaluated as
$P(x_{v},m)=\sum_{i=0}^{3}|\langle i,x_{v}|\mathcal{U}_{b_{0}}\mathcal{U}_{b_{1}}\cdots\mathcal{U}_{b_{s}}|\psi_{in}\rangle|^{2}\,$, (12)
To generate the hash function, we choose an initial state of the form
$|\psi_{in}\rangle=\frac{1}{2}\sum_{i=0}^{3}|i\rangle\otimes|x_{v}=0\rangle\,$, (13)
and perform the following steps:
  1. 1.
    Choose the values of the set of parameters $(N_{v},l_{00},\tilde{l}_{00},l_{11},\tilde{l}_{11})$. For the experiment performed in this paper we choose $(16,0.01,1.0,0.1,0.01)$.
  2. 2.
    Run quantum walk on Hanoi network HN4 as described above and in the previous section.
  3. 3.
    Measure the the probability distribution $P(x_{v},m)$ as the output.
  4. 4.
    Post-processing: Multiply probability amplitude $\sqrt{P(x_{v},m)}$ by $10^{l}$. Obtain $h_{x_{v}}=\sqrt{P(x_{v},m)}\times 10^{l}$ $mod~2^{k}$, for $10^{l}\geq 2^{k}$. The precision level, $l$, can be increased accordingly as the size of the network increases, which reduces the differences between probability amplitudes, requiring more decimal places to be included. We choose $l=5,k=16$, which makes each $h_{x_{v}}$ a $16$-bit binary string. Concatenate all $h_{x_{v}}$s to obtain a $L=N_{v}\times 16$ bit hash function as $h(m)=\left(h_{0},h_{1},\cdots,h_{N_{v}-1}\right)$.
Figure 3: Plot of the hash values $h(m_{1})$, $h(m_{2})$, $h(m_{3})$, $h(m_{4})$, and $h(m_{5})$. Red stars are the bit-values and blue curves are step plots of the hash values respectively.
$N=2000$ $N=5000$ $N=10000$ Mean
$B_{min}$ $91$ $95$ $92$ $92.6667$
$B_{max}$ $147$ $148$ $154$ $149.6667$
$\bar{B}$ $122.52$ $122.6322$ $122.6559$ $122.6027$
$\bar{P}(\%)$ $47.8594$ $47.9032$ $47.9125$ $47.8917$
$\Delta B$ $8.0311$ $8.1753$ $8.1115$ $8.1071$
$\Delta P(\%)$ $3.1372$ $3.1935$ $3.1686$ $3.1664$
Table 2: Result for diffusion and confusion test.
IV Performance of hash function
In general, our proposed hash function has a bit-length $L=N_{v}\times k$, where $N_{v}$ is the number of vertices $x_{v}$ of the HN4 network and $k$ is the bit-length of each probability amplitude $h_{x_{v}}$. There is no conclusive analytical proof for the security of the hash function. In the following sub-sections we perform several standard statistical analysis [24, 25, 26, 29, 30, 31, 32, 33, 34, 35], which show that the hash function is secured and highly collision resistant.
IV.1 Sensitivity of hash values to change in message
From a random message, by making four different types of changes to the bit-string, we obtain four additional messages. These five messages generate five corresponding hash values, which are compared to show that a small change in the original message outputs huge changes in the hash value. The steps are the following:
  1. 1.
    Choose a random message, $m_{1}$.
  2. 2.
    Change a randomly chosen bit $1$ of the message $m_{1}$ to bit 0 to obtain the message, $m_{2}$.
  3. 3.
    Change a randomly chosen bit 0 of the message $m_{1}$ to bit $1$ to obtain the message, $m_{3}$.
  4. 4.
    Add a bit at a random position of the message $m_{1}$ to obtain the message, $m_{4}$.
  5. 5.
    Delete a bit at a random position of the message $m_{1}$ to obtain the message, $m_{5}$.
  6. 6.
    Generate hash values $h(m_{1})$, $h(m_{2})$, $h(m_{3})$, $h(m_{4})$, and $h(m_{5})$.
The messages and their corresponding $256$-bit hash values in hexadecimal format are presented in table 1. We also plot the hash values in fig. 3, which shows that a small change in the original message, $m_{1}$, results in huge changes in the hash value, showing high collision resistance property.
Figure 4: Uniform distribution of bits on hash space
IV.2 Diffusion and confusion properties
In this sub-section we study the diffusion and confusion property the hash function. The following steps are performed:
  1. 1.
    Choose a random message, $m_{1}$.
  2. 2.
    Change a bit at random position of the previous message to obtain another message, $m_{2}$.
  3. 3.
    Generate $256$-bit hash values, $h(m_{1}),h(m_{2})$, in binary format, corresponding to the pair of messages $m_{1},m_{2}$ respectively.
  4. 4.
    Count the total number of flipped bit(Hamming distance) for the two hash values. Let $h(m_{1})=\left(b_{1}^{1},b_{2}^{1},\cdots,b_{L}^{1}\right)$, $h(m_{2})=\left(b_{1}^{2},b_{2}^{2},\cdots,b_{L}^{2}\right)$ are the hash values in binary form. Then Hamming distance $B_{i}$, for the $i_{th}$ experiment among the total $N$ experiments is given by $B_{i}=\sum_{j=0}^{L-1}|b_{j}^{1}-\bar{b}_{j}^{2}|$.
  5. 5.
    Repeat 1.3.1.-3. NN times and calculate the following:
  6. 6.
    $B_{min}=\scriptsize{min}\left(\{B_{i}\}^{N}_{1}\right)$.
  7. 7.
    $B_{max}=\scriptsize{max}\left(\{B_{i}\}^{N}_{1}\right)$.
  8. 8.
    $\bar{B}=\left(\sum_{i=1}^{N}B_{i}\right)/N$.
  9. 9.
    $P=\bar{B}/L$, $P(\%)=P\times 100$
  10. 10.
    $\Delta B=\sqrt{\sum_{i=1}^{N}(B_{i}-\bar{B})^{2}}/(N-1)$
  11. 11.
    $\Delta P(\%)=\sqrt{\sum_{i=1}^{N}(B_{i}/L-P)^{2}/(N-1)}\times 100$
Diffusion and confusion test are performed for $N=2000,5000$ and $10000$ and results presented in table 2 are close to the ideal values.
Figure 5: Experimental (green) and theoretical (red) collision rate ($\%$) based on $N=10000$ pair of randomly chosen $1.5N_{v}$-bit length messages as function of the number of vertices of the HN4 network for $l=7,k=16$.
IV.3 Uniform distribution
In this sub-section we study the distribution of bits in the hash values and look for the statistical signature of the original message in the hash values. The following steps are performed:
  1. 1.
    Perform 1.3.1.-3. from sub-section IV.2.
  2. 2.
    Count the number of flipped bit at a location for $N=10000$ pair of samples. Let $h(m_{1})=\left(b_{1}^{1},b_{2}^{1},\cdots,b_{L}^{1}\right)$, $h(m_{2})=\left(b_{1}^{2},b_{2}^{2},\cdots,b_{L}^{2}\right)$ are the hash values in binary form. Then the number of flipped bit-pairs, $T_{i}$, at $i_{th}$ position for $N$ experiments is given by $T_{i}=\sum_{j=0}^{N-1}\delta_{j}\left(b_{i}^{1}-\bar{b_{i}^{2}}\right)$, where $\bar{b_{i}^{2}}$ is NOTNOT $b_{i}^{2}$. $\delta_{j}(x)=1$ for $x=0$, and $\delta_{j}(x)=0$ for $x\neq 0$. $T_{i}$ is plotted in fig. 4(upper panel) as a function of the bit location.
  3. 3.
    Repeat $1.$ and count the number of same bit-pairs, $\tilde{T_{i}}=\sum_{j=0}^{N-1}\delta_{j}\left(b_{i}^{1}-b_{i}^{2}\right)$, at a location for $N=10000$ pair of samples. $\tilde{T_{i}}$ is plotted in fig. 4(lower panel) as a function of the bit location.
We can also estimate the mean of flipped bit count [44] over the bit locations as
$\bar{T}=\sum_{i=1}^{L}T_{i}/L\,$, (14)
and the standard deviation
$\Delta T=\sqrt{\sum_{i=1}^{L}\left(T_{i}-\bar{T}\right)^{2}/(L-1)}\,.$ (15)
For $N=10000$ experiments, done in ideal conditions, one would expect the mean $\bar{T}$ to be $N/2=5000$. The departure of the mean from the ideal value $|\bar{T}-N/2|$ along with $\bar{T}$ and $\Delta T$ are presented in the table 3.
Hash function $\bar{T}$ $|\bar{T}-N/2|$ $\Delta T$
$256$-bit hash function 4790.92 209.08 10.78
Table 3: Result for uniform distribution property
$\omega=0$ $\omega=1$ $\omega\geq 2$
$W^{E}\left(\omega\right)$ $9995$ 55 0
$W^{T}\left(\omega\right)$ $9997$ 22 0
Table 4: Experimental and theoretical values of collision.
Ref. [24] Ref. [30] Ref. [26] Ref. [32] Ref. [33] Our scheme
Exp. $23.12$ [33] $10.18$ $6.33$ $1.95$ $1.46$ $0.05$
Theor. $9.32$ $9.32$ $0.06$ $0.30$ $0.21$ $0.02$
Table 5: Collision rate(%) of different hash function schemes.
Coin dim. Initial state Evolution operators
Ref. [24] 44 $|00\rangle_{v}\otimes(\alpha|00\rangle+\beta|01\rangle+\chi|10\rangle+\delta|11\rangle)_{c}$ $S_{xy}(I\otimes C_{0})/S_{xy}(I\otimes C_{1})$
Ref. [30] 22 $|0\rangle_{v}\otimes(\alpha|0\rangle+\beta|1\rangle)_{c}$ $S_{y}(I\otimes C_{0})S_{x}(I\otimes C_{0})/S_{y}(I\otimes C_{1})S_{x}(I\otimes C_{1})$
Ref. [26] 44 $|0\rangle_{v}\otimes(\alpha|00\rangle+\beta|01\rangle+\chi|10\rangle+\delta|11\rangle)_{c}$ $S_{xy}(I\otimes C_{0})/S_{xy}(I\otimes C_{1})$
Ref. [32] n1n-1 $|0\rangle_{v}\otimes\frac{1}{\sqrt{n-1}}\sum_{i=1}^{n-1}|i\rangle_{c}$ $S_{J}(I\otimes C_{0})/S_{J}(I\otimes C_{1})$
Ref. [33] 22 $|0\rangle_{v}\otimes(\cos\alpha|0\rangle+\sin\alpha|1\rangle)_{c}$ $S_{c}(I\otimes C_{00})/S_{c}(I\otimes C_{01})/S_{c}(I\otimes C_{10})/S_{c}(I\otimes C_{11})$
Our scheme 44 $\frac{1}{2}\sum_{i=0}^{3}|i\rangle_{c}\otimes|0\rangle_{v}$ $S_{0}(C_{0}\otimes I)/S_{1}(C_{1}\otimes I)$
Table 6: Comparison of resources of different hash function schemes.
IV.4 Collision analysis
One of the most important things to keep in mind, while designing a hash function, is that different messages are not mapped to the same hash value. This property is known as collision resistance, which is hard to prove for a hash function. However, we can study collision resistance property of our hash scheme by estimating the number of times $\omega$ there is a match between the two numbers at the same location of a pair of hash values corresponding to two different messages. The following experiment is performed:
  1. 1.
    Perform 1.2.1.-2. from subsection IV.2.
  2. 2.
    Generate $256$-bit hash values, $h(m_{1}),h(m_{2})$, in decimal format, corresponding to the pair of messages $m_{1},m_{2}$ respectively. For the quantum walk with $N_{v}$ vertices, there are $N_{v}$ decimal numbers in a hash value.
  3. 3.
    For the pair of hash values $h(m_{1})$ and $h(m_{2})$, count the number of pairs, which have same decimal numbers at same locations. Let $h(m_{1})=\left(d_{1}^{1},d_{2}^{1},\cdots,d_{N_{v}}^{1}\right)$, $h(m_{2})=\left(d_{1}^{2},d_{2}^{2},\cdots,d_{N_{v}}^{2}\right)$ are the hash values in decimal form. Then the number of same decimal pairs at same location, $\omega=\sum_{i=1}^{N_{v}}\delta\left(d_{i}^{1}-d_{i}^{2}\right)$, $\omega=0,1,2,\cdots,N_{v}$.
  4. 4.
    Repeat 1.3.1.-3. for $N=10000$ times to obtain the number of pairs $W^{E}(\omega)$ having a fixed $\omega$.
Theoretical value can be calculated assuming binomial distribution as
$W^{T}(\omega)=$ $N\times\scriptsize{prob}(\omega)=$ (16)
$N\frac{N_{v}!}{\omega!(N_{v}-\omega)!}\left(\frac{1}{2^{k}}\right)^{\omega}\left(1-\frac{1}{2^{k}}\right)^{N_{v}-\omega}\,$,
$W^{T}(\omega)$ and $W^{E}(\omega)$ presented in table 4 shows that the experimental values are close to the theoretical values. The experimental collision rare of $0.05\%$ is very close to the theoretical collision rate of $0.02\%$, indicating very strong collision resistance. Note that such high collision resistance of our hash function is attributed to the suitable choice of values for the long-range-edge weights $l_{00},\tilde{l}_{00},l_{11},\tilde{l}_{11}$. The set of values $l_{00}=0.01,\tilde{l}_{00}=1.0,l_{11}=0.1,\tilde{l}_{11}=0.01$ chosen heuristically in this article is one of the many such suitable set of values for which the experimental collision rate $W^{E}(\omega)$ is close to the theoretical collision rate $W^{T}(\omega)=0.02\%$. It is also possible to obtain optimum values for the long-range-edge weights by optimizing the collision rate of the hash scheme with respect to the corresponding theoretical collision rate.
IV.5 Resistance to birthday attack
In this paper we have discussed hash values with a bit-length of $L=N_{v}\times k=16\times 16=256$, which means we need $2^{L/2}=2^{128}$ trials to find if two hash values collide with a probability of $0.5$, which suggests high resistance to birthday attack. Moreover, we can further increase bit-length of the hash value by increasing the length of the lattice $N_{v}$. In cryptography, fixed-size classical hash functions up-to $512$-bit hash values are currently being used in practice. Our hash function can easily be extended to $512$-bit hash value, corresponding to $N_{v}=32$, and beyond with high collision resistance property by appropriately tuning the precision level. Fig. 5 shows it has good scaling property as the collision rate of the hash function, represented by the dashed green curve is very close to the theoretical collision rate, represented by the solid red curve.
V Comparison with other quantum walk based hash schemes
One of the most important indicators for a secured hash function is the collision rate. The lower the collision rate the better is the hash function. A comparison of the collision rates(experimental and theoretical) of different hash schemes is presented in table 5, where the theoretical values are obtained using the binomial distribution in eq. (16). Since the required details for calculating $W^{T}$ for ref. [24] is not available, we calculate its theoretical value of collision in table 5, assuming $N_{v}=25$ vertices and each vertex contributing $8$ bit to a $200$-bit hash value and performing collision test for $N=10000$ times. Collision rate of our hash scheme, presented in the seventh column of table 5, shows that the experimental value $0.05$ is very close to the theoretical value $0.02$. Collision rate of our scheme can further be improved using higher precision label as shown in fig. 5. A comparison of the resource requirements of our scheme along with others are presented in table 6. Two important aspects of our proposed scheme, which are responsible for such high collision resistance property, are the following: Firstly, we have introduced two additional long-range edges to the two regular edges that allows us to tune the parameters in such a way that the probability distribution can be modified. Secondly, in addition to the coin operator we also control the shift operator by the message bits as opposed to the other quantum walk based hash functions which only control the coin operator by the message.
VI Conclusion
Quantum walk can be used to generate hash function by controlling the evolution operator $\mathcal{U}=S\left(C\otimes I\right)$. In most of the works in the literature, message controls only the coin operator $C$, depending on the bit values of the message. However, in this paper we are controlling both the coin operator $C$ and shift operator $S$. For the bit value 0, we use the evolution operator $\mathcal{U}_{0}=S_{0}\left(C_{0}\otimes I\right)$ and for the bit value $1$, we use the evolution operator $\mathcal{U}_{1}=S_{1}\left(C_{1}\otimes I\right)$, where $C_{0},C_{1}$ are Grover operators with different underlying coin states controlled by the bit values and $S_{0},S_{1}$ are flip-flop and standard shift operator respectively for the HN4 network.
Our proposed quantum walk based hash function is executed on the one-dimensional periodic lattice, equipped with extra long-range edges, known as HN4 network. It allows us to obtain a hash function, which is highly resistant to the collision in comparison with other quantum walk based hash functions as can be seen from table 5 that the collision rate is very low. It is also important to note that the the security of our scheme, including any other quantum hash schemes, is guaranteed by the Holevo theorem [45], whereas, security of classical hash functions is based on computational security of a hard mathematical function, that can be challenged by quantum algorithms.
The effect of noise on the security of hash function is not considered in this article. However, applications based on quantum walk, and more generally quantum computing models are known to be susceptible to noise. Specifically, in our case, unitary noise is important, as longer messages require more iteration steps of the quantum walk, contributing to greater decoherence related effect in the final state. Note that the noise usually disrupts the precision of the quantum operations, which can affect the deterministic nature of quantum hash function and ultimately affecting security of a quantum-based hash function as suggested in ref. [46]. In this context it also should be noted that not all noises are bad, as reported in ref. [47], broken-line decoherence type noise on a cycle improves the performance of the quantum walk based hash function. Detailed investigation on the pros and cons of different types of noises on the quantum-based hash function, including quantum walk based hash function can be an interesting future research direction. One of the possible noise mitigation techniques in our scheme would be to reduce the required number of iteration steps for a given message length by allowing more than one bit of the message to control the evolution operator at a fixed iteration step. It will also be interesting to see how the long-range edges of the HN4 network affect the unwanted noise.
Data availability Statement: The data generated during and/or analyzed during the current study is included in the article.
Conflict of interest: The authors have no competing interests to declare that are relevant to the content of this article.

### v2

@@
 Based on the number of edges at each vertex, there exist two important type of networks [37]: Hanoi network of degree three (HN3) and of degree four (HN4).
-Extensions of these typical Hanoi networks with even more degrees exist, but in this article, we only consider HN4 network for our study of the hash function.
+Extensions of these typical Hanoi networks with even more degrees exist, but, only HN3 and HN4 networks are preferable, since they require the smallest number of qubits ($\mbox{qubits}=2$) to represent the edges of the one-dimensional periodic lattice with extra long-range edges, which is preferable for the implementation in NISQ devices that have limited and qubits and low connectivity among qubits.
+In this article we have considered the HN4 network because the coin operator can be easily implemented in quantum circuits since all the four basis states of the two-qubit coin space are utilized, however we have numerically checked that the quantum hash function using the HN3 network also has similar performance.
 The main structure of this network is a one-dimensional periodic lattice of $N_{v}=2^{n}$ vertices.
@@
 However, in the case of hash function generation we need to modify the evolution operator $\mathcal{U}$ in each iteration step, which will be discussed in detail in the next section.
-One of the important properties of the quantum walk on a finite one-dimensional lattice with $t$ time steps is that, for even $N$, probability distribution at location $x$ is zero [15] when $x+t$ is odd as can be seen from fig.
+One of the important properties of the quantum walk on a finite one-dimensional lattice with $t$ time steps is that, for even $N$, the probability distribution at location $x$ is zero [15] when $x+t$ is odd, as can be seen from Fig.
 2(a).
-Therefore, the probability distribution is zero at even or add vertices for odd or even $t$ respectively.
+Therefore, the probability distribution is zero at even or odd vertices for odd or even $t$ respectively.
 However, for odd $N$ the probability distribution is nonzero at all vertex locations for large enough $t$.
-So, normally, one cannot use even length one-dimensional periodic lattice for the generation of hash function, because it leads to very high collision rates due the predictable periodic zero probability distributions.
-The HN4 network, which we use in this study, allows us to consider even-length one-dimensional periodic lattice with no periodic zero-probability distribution as can be seen from fig.
+So, normally, one cannot use even-length one-dimensional periodic lattice [31] for the generation of a hash function, because it leads to very high collision rates due to the predictable periodic zero probability distributions.
+The HN4 network, which we use in this study, allows us to consider an even-length one-dimensional periodic lattice with no periodic zero-probability distribution as can be seen from Fig.
 2(b).
@@
 $|\psi_{f}\rangle=\mathcal{U}_{b_{0}}\mathcal{U}_{b_{1}}\cdots\mathcal{U}_{b_{s}}|\psi_{in}\rangle\,.$ (10)
-For example, for a message $m=`110100^{\prime}$, the final state would be $|\psi_{f}\rangle=\mathcal{U}_{0}\mathcal{U}_{0}\mathcal{U}_{1}\mathcal{U}_{0}\mathcal{U}_{1}\mathcal{U}_{1}|\psi_{in}\rangle$.
+For example, for a message $m=`110100^{\prime}$, the final state would be $|\psi_{m}\rangle=\mathcal{U}_{0}\mathcal{U}_{0}\mathcal{U}_{1}\mathcal{U}_{0}\mathcal{U}_{1}\mathcal{U}_{1}|\psi_{in}\rangle$.
 The iteration steps for the generation of our hash function, obtained from eq.
 (10) increases linearly with the bit-length $s$ of the message.
-Although, the results presented in this article are based on noiseless analysis, in real-life scenarios, a long bit-length can transform a quantum probability distribution of the quantum walk into a classical Gaussian-like distribution in the presence of unitary noise [41].
+Although, the results presented in this article are based on noiseless analysis, in real-life scenarios, a longer bit-length, corresponding to more iteration steps, can transform a quantum probability distribution of the quantum walk into a classical Gaussian-like distribution in the presence of unitary noise [41].
+Noise associated with more iterations steps significantly degrades the probability amplitudes in Grover search algorithms also, which can be mitigated with strategies involving less iteration steps [42].
 Correction methods such as error correction [43], and decoherence free sub-spaces [44] require significant overhead, which is a luxury in the current NISQ era.
@@
 [33], where each evolution operator is controlled by two bits of the message instead of one in eq.
-(10), without compromising the performance our hash function.
+(10), without compromising the performance of our hash function.
 The final state for the hash function then becomes
@@
 [33].
+We acknowledge that a comprehensive performance validation under realistic noisy conditions remains an essential issue.
+Future work will involve detailed simulations to assess the impact of varying noise levels on collision rates, distribution uniformity, and the effectiveness of our proposed scheme.
 The final probability for the walker to be at $x_{v}$ location can be evaluated as
@@
 There is no conclusive analytical proof for the security of the hash function.
+However, the process of our quantum walk based hash function involves measurement of final quantum state in computational basis and in the post-processing stage module $2^{k}$ operation, both of which being many-to-one mappings, are irreversible one-way processes.
+Therefore, it is resistant to preimage attack, as given a hash value $y$ it is hard to find its corresponding message $m$, such that $H(m)=y$.
+It is also resistant to second preimage attacks, since given a message $m_{1}$ it is computationally infeasible because of the irreversible nature of the involved processes, to find another message $m_{1}$ such that $H(m_{1})=H(m_{2})$.
 In the following sub-sections we perform several standard statistical analysis [24, 25, 26, 29, 30, 31, 32, 33, 34, 35], which show that the hash function is secured and highly collision resistant.
@@
 Diffusion and confusion test are performed for $N=2000,5000$ and $10000$ and results presented in table 2 are close to the ideal values.
-Figure 5: Experimental (green) and theoretical (red) collision rate ($\%$) based on $N=10000$ pair of randomly chosen $1.5N_{v}$-bit length messages as function of the number of vertices of the HN4 network for $l=7,k=16$.
+Figure 5: Experimental (green) and theoretical (red) collision rate based on $N=10000$ pair of randomly chosen $1.5N_{v}$-bit length messages as function of the number of vertices of the HN4 network for $l=7,k=16$.
 IV.3 Uniform distribution
@@
 The departure of the mean from the ideal value $|\bar{T}-N/2|$ along with $\bar{T}$ and $\Delta T$ are presented in the table 3.
+Figure 6: Experimental (solid green) and theoretical (dashed green) collision rates as function of $N=10000$ pair of randomly chosen messages of varying bit-length with HN4 network for $l=6,k=16$.
 Hash function $\bar{T}$ $|\bar{T}-N/2|$ $\Delta T$
@@
 Table 6: Comparison of resources of different hash function schemes.
+Figure 7: Comparison of collision rates for five randomly chosen set of long-range weights $\lambda_{00},\tilde{\lambda}_{00},\lambda_{11},\tilde{\lambda}_{11},$ of the HN4 network.
 IV.4 Collision analysis
@@
 $W^{T}(\omega)$ and $W^{E}(\omega)$ presented in table 4 shows that the experimental values are close to the theoretical values.
-The experimental collision rare of $0.05\%$ is very close to the theoretical collision rate of $0.02\%$, indicating very strong collision resistance.
+The experimental collision rare of $c_{E}=W^{E}(\omega)\times 100=0.05\%$ is very close to the theoretical collision rate of $c_{T}=W^{T}(\omega)\times 100=0.02\%$, indicating very strong collision resistance.
 Note that such high collision resistance of our hash function is attributed to the suitable choice of values for the long-range-edge weights $l_{00},\tilde{l}_{00},l_{11},\tilde{l}_{11}$.
-The set of values $l_{00}=0.01,\tilde{l}_{00}=1.0,l_{11}=0.1,\tilde{l}_{11}=0.01$ chosen heuristically in this article is one of the many such suitable set of values for which the experimental collision rate $W^{E}(\omega)$ is close to the theoretical collision rate $W^{T}(\omega)=0.02\%$.
-It is also possible to obtain optimum values for the long-range-edge weights by optimizing the collision rate of the hash scheme with respect to the corresponding theoretical collision rate.
+The set of values $l_{00}=0.01,\tilde{l}_{00}=1.0,l_{11}=0.1,\tilde{l}_{11}=0.01$ chosen heuristically in this article is one of the many such suitable set of values for which the experimental collision rate $c_{E}$ is close to the theoretical collision rate $c_{T}=0.02\%$.
+The only requirement is that these parameters cannot be zero, since for zero long-range weights HN4 network reduces to a cycle, which has $100\%$ collision rate.
+In Fig.
+7 we evaluated collision rates for five randomly chosen sets of values of the normalized long-range weights $(\lambda_{00}=\mathcal{N}_{0}\sqrt{l_{00}},\tilde{\lambda}_{00}=\mathcal{N}_{0}\sqrt{\tilde{l}_{00}},\lambda_{11}=\mathcal{N}_{1}\sqrt{l_{11}},\tilde{\lambda}_{11}=\mathcal{N}_{1}\sqrt{\tilde{l}_{11}})$ and found that the collision rates are very close to the theoretical value $c_{T}=0.02$ despite wide variations in the parameters, implying any random choices of the long-range weights are suitable for the hash function.
+We also evaluated trace distance $D_{m1,m2}=\sqrt{1-|\langle\psi_{m1}|\psi_{m2}\rangle|^{2}}$ between the quantum output states $|\psi_{m1}\rangle,|\psi_{m2}\rangle$ corresponding to $N=10000$ randomly chosen pair of messages $(m1,m2)$ and found that trace distance satisfies $D_{m1,m2}\geq 1-\epsilon$ with $\epsilon\leq 0.25$, suggesting strong separability between the output quantum states.
 IV.5 Resistance to birthday attack
@@
 In cryptography, fixed-size classical hash functions up-to $512$-bit hash values are currently being used in practice.
-Our hash function can easily be extended to $512$-bit hash value, corresponding to $N_{v}=32$, and beyond with high collision resistance property by appropriately tuning the precision level.
-Fig.
-5 shows it has good scaling property as the collision rate of the hash function, represented by the dashed green curve is very close to the theoretical collision rate, represented by the solid red curve.
+Our hash function can easily be extended to $512$-bit hash value, corresponding to $N_{v}=32$, and beyond with high collision resistance property by appropriately tuning the precision level as can be seen from Fig.
+5.
+It has good scaling property as the collision rate of the hash function, represented by the dashed green curve is very close to the theoretical collision rate, represented by the solid red curve.
+The parameters $l_{00},\tilde{l}_{00},l_{11},\tilde{l}_{11}$ remain same as the previous experiments in the analysis of Fig.
+5, while the numbers of vertices $N_{v}=8,16,32,64,128,256,512$ are considered for the evaluation.
+Figure 8: Comparison of collision rates for five randomly chosen set of amplitudes $a_{1},a_{2},a_{3},a_{4}$ of the initial coin state.
 V Comparison with other quantum walk based hash schemes
@@
 A comparison of the resource requirements of our scheme along with others are presented in table 6.
-Two important aspects of our proposed scheme, which are responsible for such high collision resistance property, are the following: Firstly, we have introduced two additional long-range edges to the two regular edges that allows us to tune the parameters in such a way that the probability distribution can be modified.
+Another advantage of our scheme is that, it works well for short messages also, as demonstrated in Fig.
+6.
+The experimental collision rate, represented by the dashed green curve for $N_{v}=16$ and message bit-length 4204-20, remains close to the theoretical collision rate represented by the solid green line.
+Although, for very short message bit-length <4<4, the collision rate increases significantly, because the probability does not have enough time to spread.
+On the other hand since there is predictable zero probability amplitudes for quantum walk on even-length one-dimensional lattice as already discussed in Section II, which results in $100\%$ collision rate for the hash function.
+This is the reason most of the quantum walk based hash schemes based on one-dimensional periodic lattices cannot use even-length lattice at all.
+Also they cannot use message bit-length $N_{v}$ for odd-length one-dimensional periodic lattice as the collision rate is very high.
+Although the initial state for the analysis is chosen to be equal superposition of all the coin basis states at the origin of the vertex, i.e., $|\psi_{in}\rangle=\frac{1}{2}\sum_{i=0}^{3}|i\rangle\otimes|x_{v}=0\rangle$ for simplicity, however, we can randomly choose the amplitudes $a_{1},a_{2},a_{3},a_{4}$ of the initial state $|\psi_{in}\rangle=\left(a_{1}|0\rangle+a_{2}|1\rangle+a_{3}|2\rangle+a_{4}|3\rangle\right)\otimes|x_{v}=0\rangle$ to carry out our analysis without compromising performance.
+In Fig.
+8 we evaluated collision rates for five randomly chosen sets of amplitudes and found that the collision rates are very close to the theoretical value $c_{T}=0.02$ despite wide variations in the amplitudes, implying any random choices of the initial state are suitable for the hash function.
+Two important aspects of our proposed scheme, which are responsible for such high collision resistance property, are the following: Firstly, we have introduced two additional long-range edges to the two regular edges that allow us to tune the parameters in such a way that the probability distribution can be modified.
 Secondly, in addition to the coin operator we also control the shift operator by the message bits as opposed to the other quantum walk based hash functions which only control the coin operator by the message.